OpenStudy (anonymous):

Solve.x over 4 = x minus 2 over 3 A. x = 4 B. x = 8 C. x = 2

OpenStudy (anonymous):

(x/4) = (x-2)/3 Multiply both sides by 12 3x = 4x - 8 solve for x x = 8 So, B.